OpenStudy (anonymous):

If the circumference of a circle is 50 pi , calculate the area of the circle.

myininaya (myininaya):

C=2pi*r=50*pi so 2r=50=>r=25 A=pi*r^2=pi*(25)^2

OpenStudy (anonymous):

thanx!